Find the sum of:   (c) – 312, 39 and 192  
eufgoeabhuiioQ;WQ8FHYDJ NBSCS

Grade:12

1 Answers

Harshit Singh
3 years ago

Dear Student

(-312) + (+39) + (+192) = (-231) + (-81) + (+39) + (+192)

= (-231) + (-81) + (+231)
= (-231) + (+231) + (-81)
= 0 + (-81) [(-a) + (+a) = 0] = -81


Thanks